# Beware of the Classical Benchmark Instances for the Traveling Salesman Problem with Time Windows

### v1

@@ Abstract
We propose a simple and exact informed search method for the Traveling Salesman Problem with Time Windows and Makespan objective (TSPTW-M) that solves all instances of the classical benchmark with $50$ or more customers in less than ten seconds each. Applying this algorithm as an off-the-shelf method, we also solve all but one of these instances for the Duration objective. Our main conclusion is that these instances should no longer be employed for evaluating the TSPTW-M and its Duration variant: they can be “hacked” to yield results that seem outstanding at first sight. Additionally, caution is advised when designing hard training sets for machine learning algorithms.
@@ 1 Introduction
In a Traveling Salesman Problem with Time Windows (TSPTW), a vehicle must visit a set of customers within their predefined time windows. The vehicle cannot arrive late to a customer, although it can arrive early and wait until the beginning of its time window. The journey begins at a depot to which the vehicle must return after visiting all the customers. Several variants of the TSPTW are defined, depending on the objective function. Three of the most studied variants involve minimizing the time to complete the route (TSPTW-M), the duration of the route (TSPTW-D), or the travel time of the vehicle, disregarding waiting times (TSPTW-TT). From here on, we use the acronym TSPTW to refer to one of these three variants.
@@
Research on the TSPTW began over forty years ago with foundational papers, such as ChristofidesMingozziToth1981 and Baker1983 for the TSPTW-M, Savelsbergh1985 for the TSPTW-TT, and Savelsbergh1992 for the TSPTW-D. Since then, numerous algorithms employing various solving strategies have been proposed for these three problem variants.
@@
For exact approaches, significant efforts have focused on developing strong formulations suited for branch-and-bound and branch-and-cut methods (e.g. Baker1983; LangevinDesrochersDesrosiersGelinasSoumis1993; AscheuerFischettiGroetschel2001; KaraKocAltiparmakDengiz2013), including time-expanded networks (e.g. DashGuenluekLodiTramontani2012; BolandHewittVuSavelsbergh2017), dynamic programming (both pure and incomplete) (e.g. DumasDesrosiersGelinasSolomon1995; MingozziBiancoRicciardelli1997; BalasSimonetti2001), and anytime informed search methods (e.g. FontaineDibangoyeSolnon2023; KuroiwaBeck2023). Other methods include dynamic programming within column generation and Lagrangian relaxation (e.g. BaldacciMingozziRoberti2012; TilkIrnich2017; Lera-RomeroMirandaBrontSoulignac2022), constraint programming (e.g. PesantGendreauPotvinRousseau1998; FocacciLodiMilano2002), and more recently, (multivalued) decision diagrams (e.g. GillardCoppeSchausCire2021; RudichCappartRousseau2023; CoppeGillardSchaus2024).
@@
In terms of heuristic and metaheuristic approaches, in addition to the pioneering works by Savelsbergh1985 and Savelsbergh1992, notable contributions include heuristic methods by GendreauHertzLaporteStan1998; Calvo2000; Helsgaun2017, tabu search by CarltonBarnes1996, compressed annealing by OhlmannThomas2007, and ant colony optimization (ACO) along with its Beam-ACO extensions by FavarettoMorettiPellegrini2006; Lopez-IbanezBlum2010; Lopez-IbanezBlumOhlmannThomas2013. Other significant approaches include general variable neighborhood search (e.g. daSilvaUrrutia2010; MladenovicTodosijevicUrosevic2013; AmgharCordeauGendron2019; YeBartoliniSchneider2024), iterated maximum large neighborhood search by Pralet2023, and more recently, machine learning-based solvers (e.g. CappartMoisanRousseauPremont-SchwarzCire2021; KoolHoofGromichoWelling2022; ZhengHeZhouJinLi2023; BiMaZhouSongCaoWuZhang2024; ChenGongChenLiuWangYuZhang2025; li2025lmasklearnsolveconstrained), among others.
@@
Despite the many techniques developed for the TSPTW, the performance of most algorithms is evaluated on a subset of the classical benchmark instances compiled by López-Ibáñez and Blum in https://lopez-ibanez.eu/tsptw-instances, regardless of the objective function under consideration. The common reasoning is that the validity of an instance depends more on the space of feasible routes than on the objective function. Many machine learning algorithms replicate the generation process of some of these instances to create the large datasets of “hard” instances needed for training purposes (see e.g. CappartMoisanRousseauPremont-SchwarzCire2021; KoolHoofGromichoWelling2022; BiMaZhouSongCaoWuZhang2024; li2025lmasklearnsolveconstrained). A recent exception to this rule for evaluating the TSPTW is Fontaine2023, who designs hard instances with few customers based on a benchmark by RifkiChiabautSolnon2020.
@@
From the perspective of exact solvers, the classical benchmark is nearing the end of its useful life. BaldacciMingozziRoberti2012 solved all but one of the instances for the TSPTW-TT in at most one hour; RudichCappartRousseau2023; FontaineDibangoyeSolnon2023 jointly solved all but five instances for the TSPTW-M in at most one hour; and Lera-RomeroMirandaBrontSoulignac2022 solved all but fourteen of the considered instances for the TSPTW-D in at most three hours. A natural question then arises: how will we develop the next generation of benchmarking instances? daSilvaUrrutia2010 addressed this question by creating a benchmark of large instances with $400$ customers, using similar methods to those employed for the classical instances. Interestingly, FontaineDibangoyeSolnon2023 solved the benchmark by daSilvaUrrutia2010 in a few minutes, even though it is unable to solve some instances by Fontaine2023; RifkiChiabautSolnon2020 with $30$ customers within an hour.
@@
The idea that the hardest instances for the TSPTW are those with loose time windows is widely accepted in the literature (see e.g. DumasDesrosiersGelinasSolomon1995). What sets the benchmark proposed by Fontaine2023 apart is its use of a parameter, $\beta$, to systematically widen the time windows, as suggested by AriglianoGhianiGriecoGuerrieroPlana2019 for time-dependent problems. In a recent study, RifkiSolnon2025 investigate the effects of $\beta$ and another parameter, $\alpha$, on the feasibility and hardness of Euclidean instances for the TSPTW-M. Their main finding regarding feasibility is the presence of a pronounced phase transition between the infeasible and feasible regions. On the topic of hardness, they empirically demonstrate that the exact method proposed by FontaineDibangoyeSolnon2023 benefits from tight time windows. As the authors note, their analysis is crucial for the development of next-generation of benchmark instances and for training machine learning algorithms that requires datasets with varying levels of difficulty. More recently, ChenGongChenLiuWangYuZhang2025 have questioned the time window generation processes used in previous machine learning studies, suggesting that they result in weak instances that are too easy to solve.
@@ 1.1 Our contributions
The main purpose of this article is to reinforce the conclusions of RifkiSolnon2025 and ChenGongChenLiuWangYuZhang2025, demonstrating that the classical benchmark instances with $50$ or more customers, as well as the benchmark instances by daSilvaUrrutia2010, should no longer be used to evaluate the TSPTW-M. This is because these instances can be solved by a simple exact algorithm in less than ten seconds each. Indeed, any method that includes a variant of our solver as a preprocessing step could yield misleading results that appear outstanding at first glance, regardless of how well does the algorithm performs after this preprocessing phase. Our method works by repeatedly running a simple informed search in a backward direction, prioritizing partial routes that reach the end depot earlier. Despite its simplicity, it is surprising that the algorithm can solve all the large instances, including the four that remained open in previous studies. However, the algorithm has serious limitations: it fails to solve some classical benchmark instances with fewer than $50$ customers and cannot solve any instance in the benchmark by Fontaine2023; RifkiChiabautSolnon2020 with 30 or 40 customers and loose time windows.
@@
Regarding the TSPTW-D, we show that using our solver for the TSPTW-M as an off-the-shelf method is sufficient to solve all but one of the classical benchmark instances with $50$ or more customers, including the fourteen instances that remained unsolved in the study by Lera-RomeroMirandaBrontSoulignac2022. This time, all instances are solved in less than $30$ minutes each. These results further reinforce the notion that the classical benchmark instances are particularly easy for our informed search method and, therefore, they should not be used to evaluate the efficiency of algorithms for the TSPTW-D either.
@@
Although our main conclusion is that the classical benchmark instances should be avoided when evaluating the TSPTW-M and the TSPTW-D, we believe that caution is also required when evaluating the TSPTW-TT. Moreover, we argue that our results should be considered when designing new benchmark sets or training datasets for machine learning methods. Neglecting to do so could lead to biased conclusions about the efficiency and effectiveness of the solvers (ChenGongChenLiuWangYuZhang2025). The ideas proposed by AriglianoGhianiGriecoGuerrieroPlana2019, Fontaine2023, and RifkiSolnon2025, which use a time window tightness parameter $\beta$ to widen the time windows, provide a solid foundation for generating harder benchmark instances–even with fewer customers.
@@
The fact that changing $\beta$ yields harder instances certainly helps researchers evaluate the efficiency of their methods, but it does not imply that the synthetic time windows generated are significant for real-life situations. For example, Ascheuer1996 discuss a real-life problem where tight time windows are designed to ensure that jobs are completed promptly in an online setting. The short question, whose answer remains unclear to us, is whether the classical benchmark instances are truly representative of real-world problems. An affirmative answer could suggest that our algorithm, while too simplistic as a general purpose approach, might still be valid as a preprocessing step for filtering out easy instances before applying more sophisticated techniques.
@@ 2 Problem Statement
A partial route is a nonempty sequence of vertices in $D$. A partial route $R=\langle v_{0},\ldots v_{k}\rangle$ is a route if $v_{0}=0$ and $v_{k}=k=n+1$, while it is elementary when $v_{i}\neq v_{j}$ for $\in[\![k]\!], iji\neq j$. Thus, $R$ is an elementary route if and only if it visits each customer in $D$ exactly once.
@@
The partial route $R$ is feasible for time $t$ if $t\geq a(v_{0})$ and $\delta(R,t,i)\leq b(v_{i})$ for every $i\in[\![k]\!]$. Let $\delta(R,t)=\delta(R,t,k)$ if $R$ is feasible, and $\delta(R,t)=\infty$ otherwise. We refer to $\delta(R,t)$ as the (earliest feasible) arrival time of $R$ departing at time $t$. The makespan of $R$ is $\delta(R)=\delta(R,a(v_{0}))$, while its duration is $\Delta(R)=\min\{\delta(R,t)-t\mid t\in[T]\}$. The goal of the TSPTW-M (resp. TSPTW-D) is to find an elementary route with minimum makespan (resp. duration). Observe that the makespan (resp. duration) of such optimum routes is infinity when no elementary route is feasible for time 0.
@@ 2.1 Benchmark instances
In this article, we consider $10$ sets of benchmark instances, comprising a total of $1337$ instances. The first $7$ sets contain $467$ instances and correspond to the classical instances compiled by López-Ibáñez and Blum, available at https://lopez-ibanez.eu/tsptw-instances. As explained in Section 1, the efficiency of most algorithms solving TSPTWs is evaluated on a subset of these instances. Following Fontaine2023, we refer to the individual benchmarks as Asc (Ascheuer1996), Dum (DumasDesrosiersGelinasSolomon1995), Gen (GendreauHertzLaporteStan1998), Lan (LangevinDesrochersDesrosiersGelinasSoumis1993), Ohl (OhlmannThomas2007), Pes (PesantGendreauPotvinRousseau1998), and Pot (PotvinBengio1996). The travel times in Ohl are obtained by rounding the original distances to the nearest integer. The eighth benchmark, called Ohl-D, contains the same $25$ instances as Ohl but with travel times rounded to the to the nearest tenth. This benchmark was considered in the context of the TSPTW-D. The ninth set, called DaS, comprises the $125$ instances from daSilvaUrrutia2010, while the last set, called Rif, contains the remaining $720$ instances from Fontaine2023, based on a benchmark by RifkiChiabautSolnon2020.
To provide some context for the results discussed in the subsequent sections, we briefly review the origins of these benchmark sets. Our primary focus is on the time windows, as their distributions appear to be more significant than the number of customers (see e.g. RifkiSolnon2025). We also note that many training sets for machine learning methods are generated using procedures similar to those employed in the creation of Lan, Dum, Gen, Ohl, and DaS (e.g. CappartMoisanRousseauPremont-SchwarzCire2021; KoolHoofGromichoWelling2022; BiMaZhouSongCaoWuZhang2024; li2025lmasklearnsolveconstrained).
@@
  • Dum instances were created using the same procedure as Lan, but with $\sigma=50$ for $n\in\{20,40,60\}$ and $\omega\in\{20,40,60,80,100\}$; $n=80$ and $\omega\in\{20,40,60,80\}$; $n\in\{100,150\}$ and $\omega\in\{20,40,60\}$; and $n=200$ and $\omega\in\{20,40\}$. Quoting DumasDesrosiersGelinasSolomon1995, “for narrow widths, [our algorithm] is less than exponential […,] a 250-node problem with $\ell=20$ is easy to solve in less than 10 seconds.”
@@
  • DaS was designed to “overcome the limitations of the other test sets” as daSilvaUrrutia2010 needed “instances with more than 200 customers and wider time windows to benchmark and compare algorithms.” Additionally, they believed that the method by DumasDesrosiersGelinasSolomon1995 was “a biased way to generate instances since it does not reflect the real cases. Moreover, there is no challenge to build a feasible solution because the second-nearest neighbor tour already is a feasible solution.” To address these concerns, they generate random routes to assign the time windows, with $n\in\{200,250,300,350,400\}$, $\sigma=100$, and $\omega\in\{100,200,300,400,500\}$.
@@
  • RifkiChiabautSolnon2020 designed a benchmark set for time-dependent routing problems with travel times computed from shortest paths in the road network of Lyon, using realistic traffic simulations built from real-world data. To create Rif, Fontaine2023 removed the time dependency from the instances with $n\in\{20,30,40\}$. Regarding time windows, Fontaine2023 follows the model presented by AriglianoGhianiGriecoGuerrieroPlana2019 for time-dependent instances, setting $a(v_{i})=(\beta\delta(R,0,i)-40)$ and $b(v_{i})=\delta(R,0,i)+40$ for a random route $R=\langle v_{1},\ldots,v_{n+1}\rangle$ and $\beta\in\{0,0.25,0.5,1\}$. Here, $\beta\in[1]$ represents the tightness of the time windows, from the loosest 0 to the tightest $1$.

### v2

# © 2026. This manuscript version is made available under the CC-BY-NC-ND 4.0 licence https://creativecommons.org/licences/by-nc-nd/4.0/

@@ Abstract
We propose a simple and exact method for the Traveling Salesman Problem with Time Windows and Makespan objective (TSPTW-M) that solves all instances of the classical benchmark with $50$ or more customers in less than ten seconds each. Applying this algorithm as an off-the-shelf method, we also solve all but one of these instances for the Duration objective. Our main conclusion is that these instances alone are no longer representative for evaluating the TSPTW-M and its Duration variant: their structure can be exploited to yield results that seem outstanding at first glance. Additionally, caution is advised when designing hard training sets for machine learning algorithms.
@@ 1 Introduction
In a Traveling Salesman Problem with Time Windows (TSPTW), a vehicle must visit a set of customers within their predefined time windows. The vehicle cannot arrive late to a customer, although it can arrive early and wait until the beginning of its time window. The journey begins at a depot to which the vehicle must return after visiting all the customers. Several variants of the TSPTW are defined, depending on the objective function. Three of the most studied variants involve minimizing the following: the completion time of the route, i.e., the time at which the vehicle returns to the depot (TSPTW-M; the total duration of the route, defined as the time elapsed between departure from and return to the depot (TSPTW-D); and the total travel time, which accounts only for time spent traveling between locations and disregards waiting times (TSPTW-TT). From here on, we use the acronym TSPTW to refer to one of these three variants.
@@
Research on the TSPTW began over forty years ago with foundational papers, such as Christofides et al. (1981) and Baker (1983) for the TSPTW-M, Savelsbergh (1985) for the TSPTW-TT, and Savelsbergh (1992) for the TSPTW-D. Since then, numerous algorithms employing various solving strategies have been proposed for these three problem variants.
@@
For exact approaches, significant efforts have focused on developing formulations suited for branch-and-bound and branch-and-cut methods (e.g. Baker, 1983; Langevin et al., 1993; Ascheuer et al., 2001; Kara et al., 2013), including time-expanded networks (e.g. Dash et al., 2012; Boland et al., 2017), dynamic programming (both pure and incomplete) (e.g. Dumas et al., 1995; Mingozzi et al., 1997; Balas and Simonetti, 2001), and anytime informed search methods (e.g. Fontaine et al., 2023; Kuroiwa and Beck, 2023). Other methods include dynamic programming within column generation and Lagrangian relaxation (e.g. Baldacci et al., 2012; Tilk and Irnich, 2017; Lera-Romero et al., 2022), constraint programming (e.g. Pesant et al., 1998; Focacci et al., 2002), and more recently, (multivalued) decision diagrams (e.g. Gillard et al., 2021; Rudich et al., 2023; Coppé et al., 2024).
@@
In terms of heuristic and metaheuristic approaches, in addition to the pioneering works by Savelsbergh (1985) and Savelsbergh (1992), notable contributions include heuristic methods by Gendreau et al. (1998); Calvo (2000); Helsgaun (2017), tabu search by Carlton and Barnes (1996), compressed annealing by Ohlmann and Thomas (2007), and ant colony optimization (ACO) along with its Beam-ACO extensions by Favaretto et al. (2006); López-Ibáñez and Blum (2010); López-Ibáñez et al. (2013). Other significant approaches include general variable neighborhood search (e.g. da Silva and Urrutia, 2010; Mladenović et al., 2013; Amghar et al., 2019; Ye et al., 2024), iterated maximum large neighborhood search by Pralet (2023), and more recently, machine learning-based solvers (e.g. Cappart et al., 2021; Kool et al., 2022; Zheng et al., 2023; Bi et al., 2024; Chen et al., 2025; Li et al., 2025), among others.
@@
Despite the many techniques developed for the TSPTW, the performance of most algorithms is evaluated on a subset of the classical benchmark instances compiled by López-Ibáñez and Blum (2023), regardless of the objective function under consideration. The common reasoning is that the validity of an instance depends more on the space of feasible routes than on the objective function. Many machine learning algorithms replicate the generation process of some of these instances to create the large datasets of “hard” instances needed for training purposes (see e.g. Cappart et al., 2021; Kool et al., 2022; Bi et al., 2024; Li et al., 2025). A recent exception to this rule for evaluating the TSPTW is Fontaine (2024), who designs hard instances with few customers based on a benchmark by Rifki et al. (2020).
@@
From the perspective of exact solvers, the classical benchmark is nearing the end of its useful life. Baldacci et al. (2012) solved all but one of the instances for the TSPTW-TT in at most one hour; Rudich et al. (2023); Fontaine et al. (2023) jointly solved all but five instances for the TSPTW-M in at most one hour; and Lera-Romero et al. (2022) solved all but fourteen of the considered instances for the TSPTW-D in at most three hours. A natural question then arises: how will we develop the next generation of benchmarking instances? This question was addressed by da Silva and Urrutia (2010), who created a benchmark of large instances with $400$ customers, using similar methods to those employed for the classical instances. Interestingly, Fontaine et al. (2023) solved the benchmark by da Silva and Urrutia (2010) in a few minutes, even though it is unable to solve some instances by Fontaine (2024); Rifki et al. (2020) with $30$ customers within an hour.
@@
The idea that the hardest instances for the TSPTW are those with loose time windows is widely accepted in the literature (see e.g. Dumas et al., 1995). What sets the benchmark proposed by Fontaine (2024) apart is its use of a parameter, $\beta$, to systematically widen the time windows, as suggested by Arigliano et al. (2019) for time-dependent problems. In a recent study, Rifki and Solnon (2025) investigate the effects of $\beta$ and another parameter, $\alpha$, on the feasibility and hardness of Euclidean instances for the TSPTW-M. Their main finding regarding feasibility is the presence of a pronounced phase transition between the infeasible and feasible regions. On the topic of hardness, they empirically demonstrate that the exact method proposed by Fontaine et al. (2023) benefits from tight time windows. As the authors note, their analysis is crucial for the development of next-generation of benchmark instances and for training machine learning algorithms that requires datasets with varying levels of difficulty. More recently, Chen et al. (2025) have questioned the time window generation processes used in previous machine learning studies, suggesting that they result in weak instances that are too easy to solve.
@@ 1.1 Our contributions
The main purpose of this article is to reinforce the conclusions of Rifki and Solnon (2025) and Chen et al. (2025), demonstrating that the classical benchmark instances with $50$ or more customers, as well as the benchmark instances by da Silva and Urrutia (2010), should no longer be used alone to evaluate the TSPTW-M. This is because these instances can be solved by a simple exact algorithm in less than ten seconds each. Indeed, any method that includes a variant of our solver as a preprocessing step could yield misleading results that appear outstanding at first glance, regardless of how well the algorithm performs on alternative benchmarks. Our method works by repeatedly running a simple informed search in a backward direction, prioritizing partial routes that reach the end depot earlier. Despite its simplicity, it is surprising that the algorithm can solve all the large instances, including the four that remained open in previous studies. However, the algorithm has serious limitations: it fails to solve some classical benchmark instances with fewer than $50$ customers and cannot solve any instance in the benchmark by Fontaine (2024); Rifki et al. (2020) with 30 or 40 customers and loose time windows.
@@
Regarding the TSPTW-D, we show that using our solver to minimize makespan as an off-the-shelf method is sufficient to solve all but one of the classical benchmark instances with $50$ or more customers, including the fourteen instances that remained unsolved in the study by Lera-Romero et al. (2022). This time, all instances are solved in less than $30$ minutes each. These results further reinforce the notion that the classical benchmark instances are particularly easy for our informed search method and, therefore, they should be complemented with other instances when evaluating the TSPTW-D.
@@
Although our main conclusion is restricted to the TSPTW-M and the TSPTW-D, we believe that caution is also required when evaluating the TSPTW-TT. Moreover, we argue that our results should be considered when designing new benchmark sets or training datasets for machine learning methods. Neglecting to do so could lead to biased conclusions about the efficiency and effectiveness of the solvers (Chen et al., 2025). The ideas proposed by Arigliano et al. (2019), Fontaine (2024), and Rifki and Solnon (2025), which use a time window tightness parameter $\beta$ to widen the time windows, provide a solid foundation for generating harder benchmark instances–even with fewer customers.
@@
The fact that changing $\beta$ yields harder instances certainly helps researchers evaluate the efficiency of their methods, but it does not imply that the synthetic time windows generated are significant for real-life situations. For example, Ascheuer (1996) discuss a real-life problem where tight time windows are designed to ensure that jobs are completed promptly in an online setting. The short question, whose answer remains unclear to us, is whether the classical benchmark instances are truly representative of real-world problems. An affirmative answer could suggest that our algorithm, while too simplistic as a general purpose approach, might still be valid as a preprocessing step for filtering out easy instances before applying more sophisticated techniques.
@@ 2 Problem Statement
A partial route is a nonempty sequence of distinct vertices in $D$. A partial route $R=\langle v_{0},\ldots v_{k}\rangle$ is a route if $v_{0}=0$ and $v_{k}=k=n+1$, i.e., $R$ visits each customer in $D$ exactly once.
@@
The partial route $R$ is feasible for time $t$ if $t\geq a(v_{0})$ and $\delta(R,t,i)\leq b(v_{i})$ for every $i\in[\![k]\!]$. Let $\delta(R,t)=\delta(R,t,k)$ if $R$ is feasible, and $\delta(R,t)=\infty$ otherwise. We refer to $\delta(R,t)$ as the (earliest feasible) arrival time of $R$ departing at time $t$. The makespan of $R$ is $\delta(R)=\delta(R,a(v_{0}))$, while its duration is $\Delta(R)=\min\{\delta(R,t)-t\mid t\in[T]\}$. The goal of the TSPTW-M (resp. TSPTW-D) is to find a route with minimum makespan (resp. duration). Observe that the makespan (resp. duration) of such optimum routes is infinity when no route is feasible for time 0.
@@ 2.1 Benchmark instances
In this article, we consider $10$ sets of benchmark instances, comprising a total of $1337$ instances. The first $7$ sets contain $467$ instances and correspond to the classical instances compiled by López-Ibáñez and Blum (2023). As explained in Section 1, the efficiency of most algorithms solving TSPTWs is evaluated on a subset of these instances. Following Fontaine (2024), we refer to the individual benchmarks as Asc (Ascheuer, 1996), Dum (Dumas et al., 1995), Gen (Gendreau et al., 1998), Lan (Langevin et al., 1993), Ohl (Ohlmann and Thomas, 2007), Pes (Pesant et al., 1998), and Pot (Potvin and Bengio, 1996). The travel times in Ohl are obtained by rounding the original distances to the nearest integer. The eighth benchmark, called Ohl-D, contains the same $25$ instances as Ohl but with travel times rounded to the to the nearest tenth. This benchmark was considered in the context of the TSPTW-D. The ninth set, called DaS, comprises the $125$ instances from da Silva and Urrutia (2010), while the last set, called Rif, contains the remaining $720$ instances from Fontaine (2024), based on a benchmark by Rifki et al. (2020).
To provide some context for the results discussed in the subsequent sections, we briefly review the origins of these benchmark sets. Our primary focus is on the time windows, as their distributions appear to be more significant than the number of customers (see e.g. Rifki and Solnon, 2025). We also note that many training sets for machine learning methods are generated using procedures similar to those employed in the creation of Lan, Dum, Gen, Ohl, and DaS (e.g. Cappart et al., 2021; Kool et al., 2022; Bi et al., 2024; Li et al., 2025).
@@
  • Dum instances were created using the same procedure as Lan, but with $\sigma=50$ for $n\in\{20,40,60\}$ and $\omega\in\{20,40,60,80,100\}$; $n=80$ and $\omega\in\{20,40,60,80\}$; $n\in\{100,150\}$ and $\omega\in\{20,40,60\}$; and $n=200$ and $\omega\in\{20,40\}$. Quoting Dumas et al. (1995), “for narrow widths, [our algorithm] is less than exponential […,] a 250-node problem with $\ell=20$ is easy to solve in less than 10 seconds.”
@@
  • DaS was designed to “overcome the limitations of the other test sets” as da Silva and Urrutia (2010) needed “instances with more than 200 customers and wider time windows to benchmark and compare algorithms.” Additionally, they believed that the method by Dumas et al. (1995) was “a biased way to generate instances since it does not reflect the real cases. Moreover, there is no challenge to build a feasible solution because the second-nearest neighbor tour already is a feasible solution.” To address these concerns, they generate random routes to assign the time windows, with $n\in\{200,250,300,350,400\}$, $\sigma=100$, and $\omega\in\{100,200,300,400,500\}$.
@@
  • Rifki et al. (2020) designed a benchmark set for time-dependent routing problems with travel times computed from shortest paths in the road network of Lyon, using realistic traffic simulations built from real-world data. To create Rif, Fontaine (2024) removed the time dependency from the instances with $n\in\{20,30,40\}$. Regarding time windows, Fontaine (2024) follows the model presented by Arigliano et al. (2019) for time-dependent instances, setting $a(v_{i})=(\beta\delta(R,0,i)-40)$ and $b(v_{i})=\delta(R,0,i)+40$ for a random route $R=\langle v_{1},\ldots,v_{n+1}\rangle$ and $\beta\in\{0,0.25,0.5,1\}$. Here, $\beta\in[1]$ represents the tightness of the time windows, from the loosest 0 to the tightest $1$.